# Dimitrios M. Thilikos ${}^{,}$

### v3

@@ 3.2 Growing a wall from a well-linked set
Now, consider the separation $(A\cap X,B\cup Y).$ Since $|Y\cap S|>\alpha|S|$ it follows that $|(B\cup Y)\cap S|>\alpha|S|.$ From the assumption that $F_{i}$ is $S$-free this means that $(A\cap X,B\cup Y)$ is of order at least $i+1.$ Thus, by the submodularity of separations in undirected graphs, it follows that $(A\cup X,B\cap Y)$ is of order at most $i+1.$ Sine $A$ is a proper subset of $A\cup X,$ our construction of $(A,B)$ implies that $|(A\cup X)|>\alpha|S|.$ However, this means that $A\cup X\cup(B\cap Y)=V(G)$ while $(A,B),(X,Y),(B\cap Y,A\cup X)\in\mathcal{T}_{S}.$ This is a clear vialation of the tangle axioms and therefore impossible. Hence, a separation like $(X,Y)$ cannot exist and this means that $F_{i+1}$ is, in fact, $S$-free. ∎
@@
First let us show that for every $(A,B)\in\mathcal{S}_{k}$ we have $\{(A,B),(B,A)\}\cap\mathcal{T}_{F}\neq\emptyset.$ Since $|F|=3k$ this implies that $\mathcal{M}$ contains exactly one of $(A,B)$ and $(B,A)$ for evry $(A,B)\in\mathcal{S}_{k}.$
@@
Finally, we have to show that $\mathcal{T}_{F}$ is a truncation of $\mathcal{T}_{S}.$ To see this suppose there is some $(A,B)\in\mathcal{T}_{F}$ such that $(B,A)\in\mathcal{T}_{S}.$ Notice that this means that $|A\cap B|<k$ and $|B\cap F|>2k$ by definition of $\mathcal{T}_{F}.$ Hence, $|A\cap F|\leq k-1.$ Thus, $(A,B\cup F)$ is a separation of order less than $2k.$ As $F\subseteq B\cup F$ it follows from the definition of $S$-free sets that $|(B\cup F)\cap S|>\alpha|S|$ and thus, by definition of $\mathcal{T}_{S}$ we have $(A,B\cup F)\in\mathcal{T}_{S}.$ However, as $(B,A)\in\mathcal{T}_{S}$ by assumption, we now have that two small sides, namely $A$ and $B,$ of separations in $\mathcal{T}_{S}$ cover all of $V(G).$ This is clearly a contradiction to $\mathcal{T}_{S}$ being a tangle and thus we must ahve that $\mathcal{T}_{F}\subseteq\mathcal{T}_{S}.$
@@
Without loss of generality we may assume $x\in L.$ Indeed, it must now hold that $x\in L\setminus R,$ and $y\in R\setminus L$ since otherwise $(L,R)$ is witness that $X$ is not strongly linked in $G$ which contradicts our assumption. Moreover, $x$ has at least one neighbour in $L\setminus R$ as otherwise it can be moved to $R$ to form a separation of $G$ which vialoates the strong linkedness of $X.$ Similarly, $y$ has some neighbour in $R\setminus L.$
@@
Now if none of the two separations we have found is stuitable for our needs we have $|(B\cap R)\cup\{x\}\cap Y|\leq k-1$ and $|(B\cap L)\cup\{y\}|\leq k-1.$ It now follows that $|B\cap Y|\leq 2k-2.$ This, however, would imply that $|A\cap Y|\geq k+2$ which contradicts our initial assumption of $(A,B).$ This completes the proof. ∎
@@
Clearly, this procedure must stop after at most $\log(|V(G)|)$ many steps as either we reach a situation where $G[A_{i}\cup F_{i}]$ has treewidth at most $2t,$ or $A_{i}=\emptyset.$ It follows from our conditions above that in the later case we also have that $t<\mathsf{tw}(G[F_{i}\cup A_{i}])\leq 2t.$ Hence, this algorithm is always successfull. Moreover, in each step we call Subsection 3.2 along with minor checking on bounded treewidth graphs which can be done in time $2^{\mathcal{O}(t)}(k^{2})^{2t}2^{\mathcal{O}(k^{2})}\cdot|V(G)|,$ using, e.g., the minor-checking algorithm of [1]. This leads to the final running time. ∎
Notice that Subsection 3.2 can be combined with Reed’s algorithm for finding either a tree decomposition of boundd width or a $(k,2/3)$-well-linked set [38] to produce an updated version of the original result in [20, Proposition 3.6].
@@
We now partition $L_{1}\coloneqq P$ into $k$ pairwise disjoint subpaths $R_{1},\dots,R_{k}$ such that $|V(R_{i})\cap X^{\prime}|\geq 2k$ for each $i\in[k].$ For each $i\in[k]$ let $r_{i}$ be the first vertex of $R_{i}$ encountered when traversing along $L_{1}$ and let $Q_{i}$ be the unique row of $W$ containing $r_{i}.$ Finally we choose $L_{2},\dots,L_{k}$ to be $k-1$ arbitrary and distinct colums of $W$ which are all distinct from $L_{1}.$ Let $W^{\prime}$ be the $k$-subwall of $W$ with rows $Q_{1},\dots,Q_{k}$ and columns $L_{1},\dots,L_{k}.$
@@ 3.3 $\Sigma$-decompositions
We start by introducing the notions we will be using to describe “almost embeddings”’ in surfaces.
@@
Let $\Sigma$ be a surface. A $\Sigma$-decomposition of a graph $G$ is a pair $\delta=(\Gamma,\mathcal{D}),$ where $\Gamma$ is a drawing of $G$ is $\Sigma$ with crossings, and $\mathcal{D}$ is a collection of closed disks, each a subset of $\Sigma$ such that
@@
    the disks in $\mathcal{D}$ have pairwise disjoint interiors,
@@
    the boundary of each disk in $\mathcal{D}$ intersects $\Gamma$ in vertices only,
@@
    if $\Delta_{1},\Delta_{2}\in\mathcal{D}$ are distinct, then $\Delta_{1}\cap\Delta_{2}\subseteq V(\Gamma),$ and
@@
    every edge of $\Gamma$ belongs to the interior of one of the disks in $\mathcal{D}.$
Let $N$ be the set of all vertices of $\Gamma$ that do not belong to the interior of the disks in $\mathcal{D}.$ We refer to the elements of $N$ as the nodes of $\delta.$ If $\Delta\in\mathcal{D},$ then we refer to the set $\Delta-N$ as a cell of $\delta$, i.e., a cell is an closed disk where we removed at most three nodes from its boundary. We denote the set of nodes of $\delta$ by $N(\delta)$ and the set of cells by $C(\delta).$ For a cell $c\in C(\delta)$ the set of nodes that belong to the closure of $c$ is denoted by $\widetilde{c}.$ Notice that this means that the cells $c$ of $\delta$ with $\widetilde{c}\neq\emptyset$ form the edges of a hypergraph with vertex set $N(\delta)$ where $\widetilde{c}$ is the set of vertices incident with $c.$
@@
Let $G$ be a graph, $\Sigma$ be a surface and $\delta=(\Gamma,\mathcal{D})$ be a $\Sigma$-decomposition of $G.$ A cell $c\in C(\delta)$ is called a vortex if $|\widetilde{c}|\geq 4.$ Moreover, we call $\delta$ vortex-free if no cell in $C(\delta)$ is a vortex.
@@
Let $\rho=(\Gamma,\mathcal{D})$ be a rendition of $(G,\Omega)$ in a disk. If there exists a cell $c_{0}\in C(\delta)$ such that no cell in $C(\delta)\setminus\{c_{0}\}$ is a vortex we call the tripe $(\Gamma,\mathcal{D},c_{0})$ a cylindrical rendition of $(G,\Omega)$.
Let $\rho=(\Gamma,\mathcal{D})$ be a rendition of a society $(G,\Omega)$ in a surface $\Sigma.$ For every cell $c\in C(\rho)$ with $|\widetilde{c}|=2$ we select one of the components of $\mathsf{bd}(c)-\widetilde{c}.$ This selection is called a tie-breaker in $\rho$, and we assume every rendition to come equipped with a tie-breaker. Let $Q\subseteq G$ be either a cycle or a path that uses no edge of $\Sigma(c)$ for every vortex $c\in C(\rho).$ We say that $Q$ is grounded in $\rho$ if either $Q$ is a non-zero length path with both endpoints in $\pi_{\rho}(N(\rho)),$ or $Q$ is a cycle, and it uses edges of $\Sigma(c_{1})$ and $\Sigma(c_{2})$ for two distinct cells $c_{1},c_{2}\in C(\rho).$ A $2$-connected subgraph $H$ of $G$ is said to be grounded if every cycle in $H$ is grounded.
If $Q$ is grounded we define the trace of $Q$ as follows. Let $P_{1},\dots,P_{k}$ be distinct maximal subpaths of $Q$ such that $P_{i}$ is a subgraph of $\Sigma(c)$ for some cell $c$ and $Q=\bigcup_{i\in[k]}P_{k}$. Fix an index $i.$ The maximality of $P_{i}$ implies that its endpoints are $\pi(n_{1})$ and $\pi(n_{2})$ for distinct nodes $n_{1},n_{2}\in N(\rho).$ If $|\widetilde{c}|=2,$ define $L_{i}$ to be the component of $\mathsf{bd}(c)-\{n_{1},n_{2}\}$ selected by the tie-breaker, and if $|\widetilde{c}|=3,$ define $L_{i}$ to be the component of $\mathsf{bd}(c)-\{n_{1},n_{2}\}$ that is disjoint from $\widetilde{c}.$ Finally, we define $L_{i}^{\prime}$ by slightly pushing $L_{i}$ to make it disjoint from all cells in $C(\rho).$
@@
If a society $(G,\Omega)$ has depth at most $θ,$ then it has a linear decomposition of adhesion at most $2θ.$ Moreover there exists an algorithm that, given $(G,\Omega)$, either finds a transaction of order >$𝜃$ or outputs such a decomposition in time $\mathcal{O}(θ|V(G)|^{2}).$
@@
Let $\Sigma$ be a surface and $G$ be a graph. Let ${\delta=(\Gamma,\mathcal{D})}$ be a $\Sigma$-decomposition of $G.$ Every vortex $c$ defines a society ${(\Sigma(c),\Omega)},$ called the vortex society of $c,$ by saying that $\Omega$ consists of the vertices $\pi_{\delta}(n)$ for ${n\in\widetilde{c}}$ in the order given by $\Gamma.$ (There are two possible choices of $\Omega,$ namely $\Omega$ and its reversal. Either choice gives a valid vortex society.) The breadth of $\delta$ is the number of cells ${c\in C(\delta)}$ which are a vortex and the depth of $\delta$ is the maximum depth of the vortex societies ${(\Sigma(c),\Omega)}$ over all vortex cells ${c\in C(\delta)}.$
3.4 Local structure theorems
@@
  • ${\pi(N(\delta)\cap{\sf Boundary}{\Delta})\subseteq V({\sf Perimeter}{W})},$
@@
  • ${W-{\sf Perimeter}{W}}$ is a subgraph of ${\bigcup\{\Sigma(c)\mid c\subseteq\Delta\}}.$
@@
Let $\Sigma$ be a surface and ${\delta=(\Gamma,\mathcal{D})}$ be a $\Sigma$-decomposition of $G.$ We say that $\delta$ is $W$-central if there is no cell ${c\in C(\delta)}$ such that ${V(\Sigma(c))}$ includes the $W$-majority side of a separation of $G$ of order at most ${r-1}.$
More generally, let ${Z\subseteq V(G)}$ with ${\lvert Z\rvert\leq r-1},$ let Σsuperscriptnormal-$\Sigma^{\prime}$ be a surface, and let $\delta^{\prime}$ be a Σsuperscriptnormal-$\Sigma^{\prime}$-decomposition of ${G-Z}.$ Then $\delta^{\prime}$ is a $W$-central decomposition of ${G-Z}$ if for all separations ${(A,B)}$ of ${G-Z}$ of order at most ${r-|Z|-1}$ such that ${B\cup Z}$ is the majority side of the separation ${(A\cup Z,Z\cup B)}$ of $G,$ there is no cell ${c\in C(\delta^{\prime})}$ such that ${V(\Sigma_{\delta^{\prime}}(c))}$ contains $B.$

### v4

@@ 4.2 Growing a wall from a well-linked set
Now, consider the separation $(A\cap X,B\cup Y).$ Since $|Y\cap S|>\alpha|S|$ it follows that $|(B\cup Y)\cap S|>\alpha|S|.$ From the assumption that $F_{i}$ is $S$-free this means that $(A\cap X,B\cup Y)$ is of order at least $i+1.$ Thus, by the submodularity of separations in undirected graphs, it follows that $(A\cup X,B\cap Y)$ is of order at most $i+1.$ Sine $A$ is a proper subset of $A\cup X,$ our construction of $(A,B)$ implies that $|(A\cup X)|>\alpha|S|.$ However, this means that $A\cup X\cup(B\cap Y)=V(G)$ while $(A,B),(X,Y),(B\cap Y,A\cup X)\in\mathcal{T}_{S}.$ This is a clear violation of the tangle axioms and therefore impossible. Hence, a separation like $(X,Y)$ cannot exist and this means that $F_{i+1}$ is, in fact, $S$-free. ∎
@@
First let us show that for every $(A,B)\in\mathcal{S}_{k}$ we have $\{(A,B),(B,A)\}\cap\mathcal{T}_{F}\neq\emptyset.$ Since $|F|=3k$ this implies that $\mathcal{M}$ contains exactly one of $(A,B)$ and $(B,A)$ for every $(A,B)\in\mathcal{S}_{k}.$
@@
Finally, we have to show that $\mathcal{T}_{F}$ is a truncation of $\mathcal{T}_{S}.$ To see this suppose there is some $(A,B)\in\mathcal{T}_{F}$ such that $(B,A)\in\mathcal{T}_{S}.$ Notice that this means that $|A\cap B|<k$ and $|B\cap F|>2k$ by definition of $\mathcal{T}_{F}.$ Hence, $|A\cap F|\leq k-1.$ Thus, $(A,B\cup F)$ is a separation of order less than $2k.$ As $F\subseteq B\cup F$ it follows from the definition of $S$-free sets that $|(B\cup F)\cap S|>\alpha|S|$ and thus, by definition of $\mathcal{T}_{S}$ we have $(A,B\cup F)\in\mathcal{T}_{S}.$ However, as $(B,A)\in\mathcal{T}_{S}$ by assumption, we now have that two small sides, namely $A$ and $B,$ of separations in $\mathcal{T}_{S}$ cover all of $V(G).$ This is clearly a contradiction to $\mathcal{T}_{S}$ being a tangle and thus we must have that $\mathcal{T}_{F}\subseteq\mathcal{T}_{S}.$
@@
Without loss of generality we may assume $x\in L.$ Indeed, it must now hold that $x\in L\setminus R,$ and $y\in R\setminus L$ since otherwise $(L,R)$ is witness that $X$ is not strongly linked in $G$ which contradicts our assumption. Moreover, $x$ has at least one neighbour in $L\setminus R$ as otherwise it can be moved to $R$ to form a separation of $G$ which violates the strong linkedness of $X.$ Similarly, $y$ has some neighbour in $R\setminus L.$
@@
Now if none of the two separations we have found is suitable for our needs we have $|(B\cap R)\cup\{x\}\cap Y|\leq k-1$ and $|(B\cap L)\cup\{y\}|\leq k-1.$ It now follows that $|B\cap Y|\leq 2k-2.$ This, however, would imply that $|A\cap Y|\geq k+2$ which contradicts our initial assumption of $(A,B).$ This completes the proof. ∎
@@
Clearly, this procedure must stop after at most $\log(|V(G)|)$ many steps as either we reach a situation where $G[A_{i}\cup F_{i}]$ has treewidth at most $2t,$ or $A_{i}=\emptyset.$ It follows from our conditions above that in the later case we also have that $t<\mathsf{tw}(G[F_{i}\cup A_{i}])\leq 2t.$ Hence, this algorithm is always successful. Moreover, in each step we call Subsection 4.2 along with minor checking on bounded treewidth graphs which can be done in time $2^{\mathcal{O}(t)}(k^{2})^{2t}2^{\mathcal{O}(k^{2})}\cdot|V(G)|,$ using, e.g., the minor-checking algorithm of [1]. This leads to the final running time. ∎
Notice that Subsection 4.2 can be combined with Reed’s algorithm for finding either a tree decomposition of boundd width or a $(k,2/3)$-well-linked set [19] to produce an updated version of the original result in [11, Proposition 3.6].
@@
We now partition $L_{1}\coloneqq P$ into $k$ pairwise disjoint subpaths $R_{1},\dots,R_{k}$ such that $|V(R_{i})\cap X^{\prime}|\geq 2k$ for each $i\in[k].$ For each $i\in[k]$ let $r_{i}$ be the first vertex of $R_{i}$ encountered when traversing along $L_{1}$ and let $Q_{i}$ be the unique row of $W$ containing $r_{i}.$ Finally we choose $L_{2},\dots,L_{k}$ to be $k-1$ arbitrary and distinct columns of $W$ which are all distinct from $L_{1}.$ Let $W^{\prime}$ be the $k$-subwall of $W$ with rows $Q_{1},\dots,Q_{k}$ and columns $L_{1},\dots,L_{k}.$
@@ 4.3 $\Sigma$-decompositions
In this subsection we introduce the core concept that we will use to describe graphs that “almost” embed onto some surface. All of these definitions can be found in [17], also see [27] for additional explanations and illustrations. We start by introducing the notions we will be using to describe “almost embeddings” in surfaces.
@@
Let $\Sigma$ be a surface. A $\Sigma$-decomposition of a graph $G$ is a pair $\delta=(\Gamma,\mathscr{D}),$ where $\Gamma$ is a drawing of $G$ is $\Sigma$ with crossings, and $\mathscr{D}$ is a collection of closed disks, each a subset of $\Sigma$ such that
@@
    the disks in $\mathscr{D}$ have pairwise disjoint interiors,
@@
    the boundary of each disk in $\mathscr{D}$ intersects $\Gamma$ in vertices only,
@@
    if $\Delta_{1},\Delta_{2}\in\mathscr{D}$ are distinct, then $\Delta_{1}\cap\Delta_{2}\subseteq V(\Gamma),$ and
@@
    every edge of $\Gamma$ belongs to the interior of one of the disks in $\mathscr{D}.$
Let $N$ be the set of all vertices of $\Gamma$ that do not belong to the interior of the disks in $\mathscr{D}.$ We refer to the elements of $N$ as the nodes of $\delta.$ If $\Delta\in\mathscr{D},$ then we refer to the set $\Delta-N$ as a cell of $\delta,$ i.e., a cell is an closed disk where we removed at most three nodes from its boundary. We denote the set of nodes of $\delta$ by $N(\delta)$ and the set of cells by $C(\delta).$ For a cell $c\in C(\delta)$ the set of nodes that belong to the closure of $c$ is denoted by $\widetilde{c}.$ Notice that this means that the cells $c$ of $\delta$ with $\widetilde{c}\neq\emptyset$ form the edges of a hypergraph with vertex set $N(\delta)$ where $\widetilde{c}$ is the set of vertices incident with $c.$
@@
Let $G$ be a graph, $\Sigma$ be a surface and $\delta=(\Gamma,\mathscr{D})$ be a $\Sigma$-decomposition of $G.$ A cell $c\in C(\delta)$ is called a vortex if $|\widetilde{c}|\geq 4.$ Moreover, we call $\delta$ vortex-free if no cell in $C(\delta)$ is a vortex.
@@
Let $\rho=(\Gamma,\mathscr{D})$ be a rendition of $(G,\Omega)$ in a disk. If there exists a cell $c_{0}\in C(\delta)$ such that no cell in $C(\delta)\setminus\{c_{0}\}$ is a vortex we call the tripe $(\Gamma,\mathscr{D},c_{0})$ a cylindrical rendition of $(G,\Omega).$
Let $\rho=(\Gamma,\mathscr{D})$ be a rendition of a society $(G,\Omega)$ in a surface $\Sigma.$ For every cell $c\in C(\rho)$ with $|\widetilde{c}|=2$ we select one of the components of $\mathsf{bd}(c)-\widetilde{c}.$ This selection is called a tie-breaker in $\rho$, and we assume every rendition to come equipped with a tie-breaker. Let $Q\subseteq G$ be either a cycle or a path that uses no edge of $\sigma(c)$ for every vortex $c\in C(\rho).$ We say that $Q$ is grounded in $\rho$ if either $Q$ is a non-zero length path with both endpoints in $\pi_{\rho}(N(\rho)),$ or $Q$ is a cycle, and it uses edges of $\sigma(c_{1})$ and $\sigma(c_{2})$ for two distinct cells $c_{1},c_{2}\in C(\rho).$ A $2$-connected subgraph $H$ of $G$ is said to be grounded if every cycle in $H$ is grounded.
If $Q$ is grounded we define the trace [^11] Please notice that this is called the “track” in [17]. of $Q$ as follows. Let $P_{1},\dots,P_{k}$ be distinct maximal subpaths of $Q$ such that $P_{i}$ is a subgraph of $\sigma(c)$ for some cell $c$ and $Q=\bigcup_{i\in[k]}P_{k}.$ Fix an index $i.$ The maximality of $P_{i}$ implies that its endpoints are $\pi(n_{1})$ and $\pi(n_{2})$ for distinct nodes $n_{1},n_{2}\in N(\rho).$ If $|\widetilde{c}|=2,$ define $L_{i}$ to be the component of $\mathsf{bd}(c)-\{n_{1},n_{2}\}$ selected by the tie-breaker, and if $|\widetilde{c}|=3,$ define $L_{i}$ to be the component of $\mathsf{bd}(c)-\{n_{1},n_{2}\}$ that is disjoint from $\widetilde{c}.$ Finally, we define $L_{i}^{\prime}$ by slightly pushing $L_{i}$ to make it disjoint from all cells in $C(\rho).$
@@
If a society $(G,\Omega)$ has depth at most $θ,$ then it has a linear decomposition of adhesion at most $θ.$ Moreover there exists an algorithm that, given $(G,\Omega),$ either finds a transaction of order >$𝜃$ or outputs such a decomposition in time $\mathcal{O}(θ|V(G)|^{2}).$
@@
Let $\Sigma$ be a surface and $G$ be a graph. Let ${\delta=(\Gamma,\mathscr{D})}$ be a $\Sigma$-decomposition of $G.$ Every vortex $c$ defines a society ${(\sigma(c),\Omega)},$ called the vortex society of $c,$ by saying that $\Omega$ consists of the vertices $\pi_{\delta}(n)$ for ${n\in\widetilde{c}}$ in the order given by $\Gamma.$ (There are two possible choices of $\Omega,$ namely $\Omega$ and its reversal. Either choice gives a valid vortex society.) The breadth of $\delta$ is the number of cells ${c\in C(\delta)}$ which are a vortex and the depth of $\delta$ is the maximum depth of the vortex societies ${(\sigma(c),\Omega)}$ over all vortex cells ${c\in C(\delta)}.$
4.4 Structure with respect to a wall
@@
  • ${\pi(N(\delta)\cap{\sf Boundary}{\Delta})\subseteq V({\sf Perimeter}(W))},$
@@
  • ${W-{\sf Perimeter}(W)}$ is a subgraph of ${\bigcup\{\sigma(c)\mid c\subseteq\Delta\}}.$
@@
Let $\Sigma$ be a surface and ${\delta=(\Gamma,\mathscr{D})}$ be a $\Sigma$-decomposition of $G.$ We say that $\delta$ is $W$-central if there is no cell ${c\in C(\delta)}$ such that ${V(\sigma(c))}$ includes the $W$-majority side of a separation of $G$ of order at most ${r-1}.$
More generally, let ${Z\subseteq V(G)}$ with ${\lvert Z\rvert\leq r-1},$ let $\Sigma^{\prime}$ be a surface, and let $\delta^{\prime}$ be a $\Sigma^{\prime}$-decomposition of ${G-Z}.$ Then $\delta^{\prime}$ is a $W$-central decomposition of ${G-Z}$ if for all separations ${(A,B)}$ of ${G-Z}$ of order at most ${r-|Z|-1}$ such that ${B\cup Z}$ is the majority side of the separation ${(A\cup Z,Z\cup B)}$ of $G,$ there is no cell ${c\in C(\delta^{\prime})}$ such that ${V(\sigma_{\delta^{\prime}}(c))}$ contains $B.$
5 The local structure of a surface
The final piece of technology from [17] we will need is to describe how Kawarabayashi, Thomas, and Wollan capture a surface in their version of the local structure theorem [17].
The purpose of the following definitions is to introduce the concept of $\Sigma$-configurations which are meant as a strengthening of $\Sigma$-decompositions in a way that makes it possible to describe how a clique minor might appear during the inductive process of constructing the surface in the proof of Section 5 below.